Hello, my name is Sandra and I having trouble with Linear Equation Applications and I need your help with the below word problem. I think I know the answer to problems 1-3 and I will put them in for you and let me know if I am correct. Thank you...
The Smith family is taking a summer vacation in their motor home. During their trip, they maintain a constant speed of 65 mph along the highway.
The distance traveled by the Smith’s and the time t, in hours, they have traveled is given by the function D(t) = rt.

1. Which variable in the function D(t) = rt represents the constant speed the Smith’s maintain on their trip? R
2. Which variable in the function represents the range or how far the Smith’s have traveled? D
3. Which variable in the function represents the domain or how long the Smith’s have been driving? T
4. Create a specific function based on the given information. (Hint: For which variable do you have a value and can replace in the given function?)
5. Use the function you created to find the following:
D(4)=
D(8) =
D(10)=
I have to show all of the work and explain what the answers mean in the context of this problem.

Answer by mananth(16949)

1to 3 is ok
4. Create a specific function based on the given information. (Hint: For which variable do you have a value and can replace in the given function?)
t is a variable depending how much time you drive.
D(t)= 65(t)
5. Use the function you created to find the following:
D(4)= 65*4 = 260 miles
D(8) =65*8= 520
D(10)=65*10= 650
I have to show all of the work and explain what the answers mean in the context of this problem.
..
m.ananth@hotmail.ca